# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ 2.1.1 Quantile Regression Averaging
The problem addressed in this paper only makes use of different point predictors. Although this situation is not raised in the original paper (Nowotarski and Weron, 2015), the model presented there fits perfectly with the problem at hand. The idea is based on estimating the quantiles using individual point forecasts as independent variables.
Although the model is presented in the context of Day-Ahead Electricity Price Forecasting, it is perfectly generalizable to any problem. In particular, the model proposed for the quantile $\beta$ at time $t$ is
$q_{\beta}(y_{t})=\lambda_{0}(\beta)+\lambda_{1}(\beta)\hat{y}_{t,1}+\lambda_{2}(\beta)\hat{y}_{t,2}+\dots+\lambda_{m}(\beta)\hat{y}_{t,m}+\varepsilon_{t}(\beta);\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$ (1)
where $\hat{y}_{t,1},\hat{y}_{t,2},\dots,\hat{y}_{t,m}$ are the predictions of various models for the same time instant $t$. In particular, one-step ahead predictions would be obtained by:
$\hat{q}_{\beta}(y_{t})=\hat{\lambda}_{0}(\beta)+\hat{\lambda}_{1}(\beta)\hat{y}_{t,1}+\hat{\lambda}_{2}(\beta)\hat{y}_{t,2}+\dots+\hat{\lambda}_{m}(\beta)\hat{y}_{t,m}.$ (2)
To fit the models for time series problems, the use of a rolling window approach is proposed. Thus, the model described in (1) and (2) would be the particular model for one window. For another window, another estimation of the model parameters would be obtained. Figure 1 describes the process of a rolling window methodology. The window size in this procedure is chosen empirically.
@@ 2.2 Conformal Prediction
Conformal predictions were introduced in Vovk et al. (2005) to build prediction intervals (in the regression framework) that are valid with a finite number of data, without assumptions, except exchangeability, about any kind of distribution and for any predictive model.
Although the original approach, commonly referred to as Full Conformal Prediction, is not computationally feasible on a large scale, the approach known as Split Conformal Prediction (Lei et al., 2018; Papadopoulos et al., 2002) solves such problems making its use more appealing in a multitude of situations. This paper only focuses on the second approach.
Suppose we have $n$ points $(\bm{x}_{i},y_{i})\in\mathbb{R}^{m}\times\mathbb{R},\,i=1,\dots,n$ and we are interested in providing a prediction interval for the next observation $y_{n+1}$ for which $\bm{x}_{n+1}$ is known. Conformalization in its simplest form consists in making a correction to a predictor of the mean. Let $\hat{\mu}(\cdot)$ be that predictor. The steps to perform its conformalization for an objective miscoverage of $\alpha$ are:
  1. 1.
    Randomly split the $n$ known points into two sets: training, Tr, and calibration, Cal. $\text{Tr},\text{Cal}\subset\left\{1,\dots,n\right\}$.
  2. 2.
    Train the regression algorithm $\hat{\mu}$ using the data from the training set.
  3. 3.
    Compute a set of scores that determines how well the predictions of the regression algorithm fit the observations in the calibration set. This score is known as the conformity score and the most common is the absolute error of each of the points. That is,
    $\mathcal{S}=\mathcal{S}_{\text{Cal}}\cup\left\{+\infty\right\},$
    where $\mathcal{S}_{\text{Cal}}=\left\{|y_{i}-\hat{\mu}(\bm{x}_{i})|\,:i\in\text{Cal}\right\}$
  4. 4.
    Compute the $\frac{\lceil(|\mathcal{S}_{\text{Cal}}|+1)(1-\alpha)\rceil}{|\mathcal{S}_{\text{Cal}}|}$ quantile of the conformity scores. We will denote it by $Q_{1-\alpha}(\mathcal{S})$.
  5. 5.
    The final conformalized prediction interval for the observation $n+1$ is given by
    $\hat{C}_{\alpha,n+1}=\left[\hat{\mu}(\bm{x}_{n+1})-Q_{1-\alpha}(\mathcal{S}),\;\hat{\mu}(\bm{x}_{n+1})+Q_{1-\alpha}(\mathcal{S})\right]$
@@
Let $(\bm{x}_{i},y_{i})_{i=1}^{n+1}$ be exchangeable. The process of conformalizing a conditional mean predictor as described previously produces a prediction interval for the observation $n+1$, $\hat{C}_{\alpha,n+1}$, such that
$\mathbb{P}\left(y_{n+1}\in\hat{C}_{\alpha,n+1}\right)\geq 1-\alpha.$
If, in addition, the scores $\mathcal{S}_{\text{Cal}}$ have a continuous joint distribution, we also have:
$\mathbb{P}\left(y_{n+1}\in\hat{C}_{\alpha,n+1}\right)\leq 1-\alpha+\dfrac{1}{\#\text{Cal}+1}.$
@@ 2.2.1 Conformalized Quantile Regression (CQR)
Although guaranteeing this property is not possible (Vovk, 2012; Lei and Wasserman, 2014), a variety of works have been developed to approximate it as best as possible. The most popular of these is probably the Conformalized Quantile Regression (CQR) proposed in Romano et al. (2019). CQR follows the conformal methodology to correct the coverage obtained by estimating the quantiles through a quantile regression procedure. Although there is no theoretical result related to conditional coverage, as the correction is performed on these estimated conditional quantiles, it is expected that heteroscedasticity is captured with much better quality than with the traditional conformal approach.
The CQR methodology for an objective miscoverage rate of $\alpha$ could be summed up as:
  1. 1.
    Randomly split the $n$ known points into two sets: training, Tr, and calibration, Cal. $\text{Tr},\text{Cal}\subset\left\{1,\dots,n\right\}$.
  2. 2.
    Obtain a first approximation of $l_{\alpha,i}$ and $u_{\alpha,i},\,\hat{l}_{\alpha,i}$ and $\hat{u}_{\alpha,i},\,i\in\text{Cal}\cup\{n+1\}$ by some quantile regression algorithm using only the training set for training. Predict over the calibration set and the observation of interest $n+1$.
  3. 3.
    Compute the following conformity scores:
@@
    where $\mathcal{S}_{i}=\max\left\{y_{i}-\hat{u}_{\alpha,i},\hat{l}_{\alpha,i}-y_{i}\right\}\,\forall\,i\in\text{Cal}$
  4. 4.
    Compute the $\frac{\lceil(|\mathcal{S}_{\text{Cal}}|+1)(1-\alpha)\rceil}{|\mathcal{S}_{\text{Cal}}|}$ quantile of the conformity scores. We will denote it by $Q_{1-\alpha}(\mathcal{S})$.
  5. 5.
    The final conformalized prediction interval for the observation $n+1$ is given by
    $\hat{C}_{\alpha,n+1}=\left[\hat{l}_{\alpha,n+1}-Q_{1-\alpha}(\mathcal{S}),\hat{u}_{\alpha,n+1}+Q_{1-\alpha}(\mathcal{S})\right]$
@@
Let $\left(\bm{x}_{i},y_{i}\right)_{i=1}^{n+1}$ be exchangeable. Applying CQR $\left(\bm{x}_{i},y_{i}\right)_{i=1}^{n}$ produces a prediction interval $\hat{C}_{\alpha,n+1}$ such that:
$\mathbb{P}\left(y_{n+1}\in\hat{C}_{\alpha,n+1}\right)\geq 1-\alpha.$
@@
$\mathbb{P}\left(y_{n+1}\in\hat{C}_{\alpha,n+1}\right)\leq 1-\alpha+\dfrac{1}{\#\text{Cal}+1}.$
The algorithm has shown such positive qualities that it has become the standard in Conformal Prediction for regression. For more details of the CQR algorithm we refer to Romano et al. (2019).
@@ 2.2.2 Adaptive Conformal Inference (ACI)
CQR or any other conformal algorithm following the presented scheme depends on the condition of exchangeability between observations. In time series, which are the problems we are interested in, this condition is not fulfilled. Removing the condition of exchangeability while maintaining the validity property of the intervals has been one of the primary research objectives in the field. One such work is the Adaptive Conformal Inference (ACI) method proposed by Gibbs and Candès (2021).
@@
$\begin{cases}\alpha_{t+1}&=\alpha_{t}+\gamma(\alpha^{*}-\text{err}_{t})\\ \text{err}_{t}&=\begin{cases}1&\text{if }y_{t}\not\in\hat{C}_{\alpha^{*},t}\\ 0&\text{otherwise}\end{cases}\\ \hat{C}_{\alpha^{*},t+1}&=\left[\hat{l}_{\alpha^{*},t+1}-Q_{1-\alpha_{t+1}}(\mathcal{S}),\hat{u}_{\alpha^{*},t+1}+Q_{1-\alpha_{t+1}}(\mathcal{S})\right]\\ \end{cases}$
@@ 3 Our proposal
Previous research has examined the issue of providing valid and efficient prediction intervals for individual values. A comprehensive analysis of the coverage of prediction bands, as well as of the interval lengths associated, based on this individual approach reveals significant issues which were described in Section 1.The first is that the length of the intervals varies depending on the difficulty of the observation to be predicted. That is, there should be an increasing relationship between the error of the point forecasting model and the length of the proposed interval.
To our knowledge, the only work that uses point predictors to obtain prediction intervals, as in the context in which we are, is the QRA described in Section 2.1.1. However, its approach does not take this desired property into account. We are going to propose a model in the same spirit as the QRA, but taking into account this feature we are looking for.
@@ 3.1 Heteroscedastic Quantile Regression (HQR)
It is clear that having different predictors of the expected value can provide information on the safety of the prediction: in very common situations for the model, i.e. in areas where the space of regressor variables is highly explored, all forecasters are likely to obtain very similar predictions. However, in the more unfamiliar situations, which generally correspond to unexplored areas where models have to extrapolate, the forecasts start to differ and, in particular, the error of the models in such cases is generally larger (Figure 2). In other words, the uncertainty in unknown conditions is higher than in typical conditions.
@@
We believe that a good indicator of the level of exploration of the explanatory features space is a dispersion measure of the prediction of the different models. Thus, denoting by $\overline{\hat{y}}_{t}=\dfrac{1}{M}\displaystyle\sum_{i=1}^{M}\hat{y}_{t,i}$ and $s_{\hat{y}_{t}}^{2}=\dfrac{1}{M}\displaystyle\sum_{i=1}^{M}\left(\hat{y}_{t,i}-\overline{\hat{y}}_{t}\right)^{2}$ the following quantile regression model is proposed:
$q_{t}(\beta)=\lambda_{0}(\beta)+\lambda_{1}(\beta)\overline{\hat{y}}_{t}+\lambda_{2}(\beta)s_{\hat{y}_{t}}+\varepsilon_{t}(\beta),\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$ (3)
where the parameters are obtained by minimizing the pinball loss (2.1) and vary over time in the same way as in (1). In particular, for the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles of interest, we have:
$\begin{cases}\hat{q}_{T+1}(\frac{\alpha}{2})&=\hat{\lambda}_{0}(\frac{\alpha}{2})+\hat{\lambda}_{1}(\frac{\alpha}{2})\overline{\hat{y}}_{T+1}+\hat{\lambda}_{2}(\frac{\alpha}{2})s_{\hat{y}_{T+1}}\\ \hat{q}_{T+1}(1-\frac{\alpha}{2})&=\hat{\lambda}_{0}(1-\frac{\alpha}{2})+\hat{\lambda}_{1}(1-\frac{\alpha}{2})\overline{\hat{y}}_{T+1}+\hat{\lambda}_{2}(1-\frac{\alpha}{2})s_{\hat{y}_{T+1}}\\ \end{cases}$
@@
Note that (3) is actually an extension of the QRA model defined in (1). In the case of the QRA model, what is being done is to estimate the mean through a weighted average, which results in different values of the coefficients $\lambda_{1},\dots,\lambda_{M}$. That is, the QRA model is a model of the type
$q_{t}(\beta)=\lambda_{0}(\beta)+\lambda_{1}(\beta)\overline{\hat{y}}_{t}+\varepsilon_{t}(\beta),\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$
where the estimation $\overline{\hat{y}}_{t}$ is not done with equal weights. In that sense, we are extending the model with a further component that refers to a first assessment of the level of uncertainty that exists. Because this extension is directly related to the heteroscedasticity of the predictions, the model has been named Heteroscedastic Quantile Regression (HQR).

### v2

@@ 2.2.1 Quantile Regression Averaging
The core idea in Nowotarski and Weron (2015) is to estimate the quantiles by treating the individual point forecasts as independent variables. While the original work does not explicitly restrict itself to using only this information, the model presented there is well-suited to the current problem, which is to produce prediction intervals relying solely on different point predictors.
Although the model is presented in the context of Day-Ahead electricity price forecasting, it is perfectly generalizable to any regression problem. In particular, the model proposed for the quantile $\beta$ at time $t$ is
$q_{\beta}(Y_{t}|\bm{y}_{t})=\lambda_{0}(\beta)+\lambda_{1}(\beta)\widehat{y}_{t,1}+\lambda_{2}(\beta)\widehat{y}_{t,2}+\dots+\lambda_{m}(\beta)\widehat{y}_{t,M}+\varepsilon_{t}(\beta);\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$ (3)
where $Y_{t}$ is the random variable associated to instant $t$ and $\bm{y}_{t}=\left(\widehat{y}_{t,1},\widehat{y}_{t,2},\dots,\widehat{y}_{t,M}\right)$ are predictions of the mean from $M$ different models for the same time instant $t$. In particular, one-step ahead predictions would be obtained by:
$\widehat{q}_{\beta}(Y_{T+1}|\bm{y}_{T+1})=\widehat{\lambda}_{0}(\beta)+\widehat{\lambda}_{1}(\beta)\widehat{y}_{T+1,1}+\widehat{\lambda}_{2}(\beta)\widehat{y}_{T+1,2}+\dots+\widehat{\lambda}_{m}(\beta)\widehat{y}_{T+1,M}.$ (4)
To fit the models for time series problems, the use of a rolling window approach is proposed. Thus, the model described in equation (3) and (4) would be the particular model for one window. For another window, another estimation of the model parameters would be obtained. Figure 1 describes the process of a rolling window methodology. The window size in this procedure is chosen empirically.
@@ 2.3 Conformal Prediction
Conformal predictions were introduced in Vovk et al. (2005) to build prediction intervals (in the regression framework) that are valid with a finite number of data, without assumptions, except exchangeability, about any kind of distribution. It operates as a post-processing phase within an existing prediction pipeline, enabling the construction of valid prediction intervals without requiring any modifications to the existing forecasting process. Its fundamental base assures that intervals are valid regardless of the quality of the initial predictions, although their efficiency remains influenced by the accuracy of the preceding forecasting phase.
Although the original approach, commonly referred to as Full Conformal Prediction, is not computationally feasible on a large scale, the approach known as Split Conformal Prediction (SCP, Lei et al. (2018); Papadopoulos et al. (2002)) solves such problems making its use more appealing in a multitude of situations. This paper only focuses on the second approach.
Suppose we have $n$ points $(\bm{x}_{i},y_{i})\in\mathbb{R}^{m}\times\mathbb{R},\,i=1,\dots,n$ and we are interested in providing a prediction interval for the next observation $y_{n+1}$ for which $\bm{x}_{n+1}$ is known. Conformalization in its simplest form consists in making a correction to a prediction of the mean. Let $\widehat{\mu}(\cdot)$ be that predictor. The steps to perform its conformalization for an objective miscoverage of $\alpha$ are described in Algorithm 1.
Algorithm 1 Conformalized Mean Regression trough SCP
1:$\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, significance level $\alpha$, regression algorithm $\widehat{\mu}$
2:Randomly split the $n$ known points into two disjoint sets: training Tr and calibration Cal.
3:Train the regression algorithm $\widehat{\mu}$ using the data from the training set Tr.
4:Compute the conformity scores for the calibration set Cal using the absolute error:
$\mathcal{S}=\mathcal{S}_{\text{Cal}}\cup\{+\infty\},$
where
$\mathcal{S}_{\text{Cal}}=\{|y_{i}-\widehat{\mu}(\bm{x}_{i})|\,:\,i\in\text{Cal}\}.$
5:Compute the $(1-\alpha)$ quantile of the conformity scores, denoted as $Q_{1-\alpha}(\mathcal{S})$.
6:Construct the conformal prediction interval for observation $n+1$ as:
$\widehat{C}_{\alpha,n+1}=\left[\widehat{\mu}(\bm{x}_{n+1})-Q_{1-\alpha}(\mathcal{S}),\;\widehat{\mu}(\bm{x}_{n+1})+Q_{1-\alpha}(\mathcal{S})\right].$
7:$\widehat{C}_{\alpha,n+1}$ prediction interval of level $1-\alpha$ for the observation $n+1$.
@@
Let $(\bm{x}_{i},y_{i})_{i=1}^{n+1}$ be exchangeable. The process of conformalizing a conditional mean predictor as described in Algorithm 1 produces a prediction interval for the observation $n+1$, $\widehat{C}_{\alpha,n+1}$, such that
$\mathbb{P}\left(y_{n+1}\in\widehat{C}_{\alpha,n+1}\right)\geq 1-\alpha.$
If, in addition, the scores $\mathcal{S}_{\text{Cal}}$ have a continuous joint distribution, then:
$\mathbb{P}\left(y_{n+1}\in\widehat{C}_{\alpha,n+1}\right)\leq 1-\alpha+\dfrac{1}{\#\text{Cal}+1}.$
@@ 2.3.1 Conformalized Quantile Regression (CQR)
In Figure 2 the difference between marginal and conditional coverage is appreciated. Notice how in the case of conditional coverage, the prediction intervals are adjusted to the heteroscedasticity of the data as a function of $X$.
Figure 2: Difference between marginal coverage and conditional coverage in a toy dataset.
For any distribution-free approximation, not just Conformal Prediction, enforcing this property would require the intervals to be uninformative. That is, achieving this property in a practical way and without assuming any distribution is not possible (Vovk, 2012; Lei and Wasserman, 2014). Therefore, a variety of works have been developed to approximate it as best as possible. The most popular of these is probably the Conformalized Quantile Regression (CQR) procedure proposed in Romano et al. (2019). CQR follows the conformal methodology to correct the coverage obtained by estimating the quantiles through a quantile regression procedure. Since a quantile regression procedure cannot guarantee the desired coverage level in finite samples, CQR adjusts the interval bounds generated by this procedure. The intervals are enlarged or reduced if the empirical marginal coverage in a calibration set is found to be smaller or larger than the target level, respectively. Although there is no theoretical result related to conditional coverage, as the correction is performed on these estimated conditional quantiles, it is expected that heteroscedasticity is captured with much better quality than with the traditional conformal approach.
The CQR methodology for an objective miscoverage rate of $\alpha$ is describe in Algorithm 2.
Algorithm 2 Conformalized Quantile Regression
1:$\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, significance level $\alpha$, quantile regression algorithm $\mathcal{A}$
2:Randomly split the $n$ known points into two disjoint sets: training Tr and calibration Cal.
3:Train the quantile regression algorithm $\mathcal{A}$ using the data from the training set Tr and obtain a first approximation of $l_{\alpha,i}$ and $u_{\alpha,i},\,\widehat{l}_{\alpha,i}$ and $\widehat{u}_{\alpha,i},\,i\in\text{Cal}\cup\{n+1\}$
4:Compute the conformity scores $\mathcal{S}$
@@
where $\mathcal{S}_{i}=\max\left\{y_{i}-\widehat{u}_{\alpha,i},\widehat{l}_{\alpha,i}-y_{i}\right\}$
5:Compute the $(1-\alpha)$ quantile of the conformity scores, denoted as $Q_{1-\alpha}(\mathcal{S})$.
6:Construct the conformal prediction interval for observation $n+1$ as:
$\widehat{C}_{\alpha,n+1}=\left[\widehat{l}_{\alpha,n+1}-Q_{1-\alpha}(\mathcal{S}),\widehat{u}_{\alpha,n+1}+Q_{1-\alpha}(\mathcal{S})\right]$
7:$\widehat{C}_{\alpha,n+1}$ prediction interval of level $1-\alpha$ for the observation $n+1$.
@@
Let $\left(\bm{x}_{i},y_{i}\right)_{i=1}^{n+1}$ be exchangeable. Applying CQR $\left(\bm{x}_{i},y_{i}\right)_{i=1}^{n}$ produces a prediction interval $\widehat{C}_{\alpha,n+1}$ such that:
$\mathbb{P}\left(y_{n+1}\in\widehat{C}_{\alpha,n+1}\right)\geq 1-\alpha.$
@@
$\mathbb{P}\left(y_{n+1}\in\widehat{C}_{\alpha,n+1}\right)\leq 1-\alpha+\dfrac{1}{\#\text{Cal}+1}.$
Given that the computational complexity is minimal once the quantile regression model is trained, and valid intervals are ensured on finite and exchangeable samples, CQR is considered the standard approach in Conformal Prediction for regression tasks. Therefore, it will serve as the foundation for the conformalizations performed through ACI and WACI (Sections 2.3.2 and 3.2). For more details of the CQR algorithm we refer to Romano et al. (2019).
@@ 2.3.2 Adaptive Conformal Inference (ACI)
CQR or any other conformal algorithm following the presented scheme (Algorithm 2) depends on the condition of exchangeability among observations. In time series, which are the problems we are interested in, this condition is not fulfilled. Removing the condition of exchangeability while maintaining the validity property of the intervals has been one of the primary research objectives in the field. One such work is the Adaptive Conformal Inference (ACI) method proposed by Gibbs and Candès (2021).
@@
$\begin{cases}\alpha_{t+1}&=\alpha_{t}+\gamma(\alpha^{*}-\text{err}_{t})\\ \text{err}_{t}&=\begin{cases}1&\text{if }y_{t}\not\in\widehat{C}_{\alpha^{*},t}\\ 0&\text{otherwise}\end{cases}\\ \widehat{C}_{\alpha^{*},t+1}&=\left[\widehat{l}_{\alpha^{*},t+1}-Q_{1-\alpha_{t+1}}(\mathcal{S}),\widehat{u}_{\alpha^{*},t+1}+Q_{1-\alpha_{t+1}}(\mathcal{S})\right]\\ \end{cases}$
@@ 3 Our proposal
Previous research has focused on the issue of providing valid and efficient prediction intervals for individual values. A comprehensive analysis of the coverage of prediction bands, as well as of the interval lengths associated, based on this individual approach reveals significant issues which were described in Section 1.The first is that the length of the intervals varies depending on the difficulty of the observation to be predicted. That is, there should be an increasing relationship between the error of the point forecasting model and the length of the proposed interval.
One of the main works leveraging point predictors to build prediction intervals, as in the context of this study, is the QRA method described in Section 2.2.1. However, its approach does not account for the desired property of adapting to heteroscedasticity. In this paper, we propose a model inspired by QRA, but explicitly designed to incorporate the heteroscedasticity. This design enables our approach to simultaneously capture the aleatoric uncertainty associated with the event and the epistemic uncertainty of the predictors, ensuring that the estimated uncertainty reflects the underlying complexity of the prediction task.
@@ 3.1 Heteroscedastic Quantile Regression (HQR)
It is clear that having different predictors of the expected value can provide information on the safety of the prediction: in very common situations for the model, i.e., in areas where the space of regressor variables is highly explored, all forecasters are likely to obtain very similar predictions. However, in the more unfamiliar situations, which generally correspond to unexplored areas where models have to extrapolate, the forecasts start to differ, and, in particular, the error of the models in such cases is generally larger (Figure 3). This reflects higher epistemic uncertainty, as it stems from a lack of knowledge or a lack of data in some of these regions. Additionally, situations of high aleatoric uncertainty, where inherent randomness in the data-generating process dominates, can also lead to divergence in the predictions of the mean. In these cases, even with well-trained models, the variability in the predictions reflects the fundamental unpredictability of the process. In other words, both epistemic and aleatoric uncertainty contribute to the observed differences in the forecasts, which is a factor that should be taken into account when building prediction intervals.
@@
A good indicator of the level of exploration of the explanatory features space is a dispersion measure of the prediction of the different models. Suppose, following the model of Nowotarski and Weron (2015), that for each time $t$ we have $M$ different forecasts $\widehat{y}_{t,1},\widehat{y}_{t,2},\dots,\widehat{y}_{t,M}$ intended to predict $y_{t}$. Thus, denoting by $\overline{\widehat{y}}_{t}=\dfrac{1}{M}\displaystyle\sum_{i=1}^{M}\widehat{y}_{t,i}$ and $s_{\widehat{y}_{t}}^{2}=\dfrac{1}{M}\displaystyle\sum_{i=1}^{M}\left(\widehat{y}_{t,i}-\overline{\widehat{y}}_{t}\right)^{2}$ the following quantile regression model is proposed:
$q_{\beta}(Y_{t}|\bm{\widehat{y}}_{t})=\lambda_{0}(\beta)+\lambda_{1}(\beta)\overline{\widehat{y}}_{t}+\lambda_{2}(\beta)s_{\widehat{y}_{t}}+\varepsilon_{t}(\beta),\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$ (5)
where the parameters are obtained by minimizing the pinball loss (equation (2.2)) and vary over time using a rolling window procedure in the same way as in equation (3). In particular, for the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles of interest, we have:
$\begin{cases}\widehat{q}_{\frac{\alpha}{2}}(Y_{T+1}|\bm{\widehat{y}}_{T+1})&=\widehat{\lambda}_{0}(\frac{\alpha}{2})+\widehat{\lambda}_{1}(\frac{\alpha}{2})\overline{\widehat{y}}_{T+1}+\widehat{\lambda}_{2}(\frac{\alpha}{2})s_{\widehat{y}_{T+1}}\\ \widehat{q}_{1-\frac{\alpha}{2}}(Y_{T+1}|\bm{\widehat{y}}_{T+1})&=\widehat{\lambda}_{0}(1-\frac{\alpha}{2})+\widehat{\lambda}_{1}(1-\frac{\alpha}{2})\overline{\widehat{y}}_{T+1}+\widehat{\lambda}_{2}(1-\frac{\alpha}{2})s_{\widehat{y}_{T+1}}\\ \end{cases}$
@@
Note that equation (5) is actually an extension of the QRA model defined in (3). In the case of the QRA model, what is being done is to estimate the mean through a weighted average, which results in different values of the coefficients $\lambda_{1},\dots,\lambda_{M}$. That is, the QRA model is a model of the type
$q_{\beta}(Y_{t}|\bm{\widehat{y}}_{t})=\lambda_{0}(\beta)+\lambda_{1}(\beta)\overline{\widehat{y}}_{t}+\varepsilon_{t}(\beta),\;\mathbb{E}\left[\varepsilon_{t}(\beta)\right]=0,$
where the estimation $\overline{\widehat{y}}_{t}$ is not done with equal weights. In that sense, we are extending the model with a further component that refers to a first assessment of the level of uncertainty that exists. Because this extension is directly related to the heteroscedasticity of the predictions, the model has been named Heteroscedastic Quantile Regression (HQR).
To evaluate the significance of mean estimation (whether with equal weights or not) in the experimental section, the model denoted as HQR-W (Weighted Heteroscedastic Quantile Regression), whose expression is given by
$\begin{split}\widehat{q}_{\beta}(Y_{t}|\bm{y}_{t})&=\widehat{\lambda}_{0}(\beta)+\widehat{\lambda}_{1}(\beta)\widehat{y}_{t,1}+\widehat{\lambda}_{2}(\beta)\widehat{y}_{t,2}+\dots\\ &\quad+\widehat{\lambda}_{M}(\beta)\widehat{y}_{t,m}+\widehat{\lambda}_{M+1}(\beta)s_{\widehat{y}_{t}}+\varepsilon_{t}(\beta),\\ \mathbb{E}\left[\varepsilon_{t}(\beta)\right]&=0,\end{split}$ (6)
will also be considered. This assessment is particularly important because the number of variables included in the models can differ significantly depending on whether the approximation follows (5) or (6). While techniques such as L1 regularization (Uniejewski and Weron, 2021) or similar approaches could be employed to address this potential issue, it is essential to first determine whether such measures are necessary at all.